OpenStudy (anonymous):

Solve: absolute value of 2-x = 9

OpenStudy (anonymous):

I'd like the steps please because I don't really know what to do, subtract 2?

Parth (parthkohli):

\( \color{Black}{\Rightarrow |2 - x| = |x - 2| }\) can you do it now?

Parth (parthkohli):

\( \color{Black}{\Rightarrow |x| = y \Longrightarrow x = \pm y }\)

Parth (parthkohli):

use these properties.

OpenStudy (anonymous):

what properties? subtracting 2 and dividing by -1? I'm not sure exactly what I have to do

Parth (parthkohli):

\( \color{Black}{\Rightarrow |2 -x| = |x -2| }\) so, \( \color{Black}{\Rightarrow |x -2| = 9 \Longrightarrow x - 2= \pm9 }\) Here we have to different equations. \( \color{Black}{\Rightarrow x - 2 = 9 }\) and \( x - 2 = -9\)

OpenStudy (anonymous):

so you always make it a negative and a positive?

Parth (parthkohli):

Yes :)

OpenStudy (anonymous):

Thank you!

Parth (parthkohli):

yw :D